OpenStudy (anonymous):

I still have a question on a part of this: Show that the two functions grow at the same rate: sqrt(9^x + 2^x) and 3^x

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

what question

OpenStudy (anonymous):

Ok so when you get to \[\sqrt{\lim_{x \rightarrow \infty} 1+(2/9)^{x}}\] how do you prove that that goes to a constant?

OpenStudy (amorfide):

well you know the limit of 2/9 to the power of x goes to 0 because a decimal to the power of x will tend to 0 as x tends to infinite so you just get the limit of root 1 as x tends to infinite so you just get 1 as the limite therefore the limit of f(x)/g(x) as x tends to infinite is 1 meaning that the growth rate is the same I think that was clear idk

OpenStudy (anonymous):

I understand, but is there a mathematical way to prove it?

OpenStudy (anonymous):

I mean like to algebraically simplify it

OpenStudy (amorfide):

what do you mean like an actual proof?

OpenStudy (anonymous):

Well to simplify the (2/9)^x to 0

OpenStudy (anonymous):

Or do I not have to?

OpenStudy (amorfide):

http://prntscr.com/6ggqj2

OpenStudy (amorfide):

there is the working out as to why

OpenStudy (anonymous):

How did you go to the part that is above the gray part to the gray part? (if you know what I mean at all)

OpenStudy (anonymous):

How did you do these 3 lines http://prntscr.com/6ggsfr

OpenStudy (amorfide):

idk how to explain the continuity thing uhm @e.mccormick @Nnesha

OpenStudy (anonymous):

Its alright :/

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (amorfide):

if f(x) is continuous around x=g(a) and g(x) is continuous around x=a, then \[\lim_{x \rightarrow a} f(g(x)=f(\lim_{x \rightarrow a}g(x))\] compute \[e ^{\lim_{x \rightarrow infty}(xln(2/9))}\] \[\lim_{x \rightarrow \infty}(xln (\frac{ 2 }{ 9 }))=-\infty \]

OpenStudy (amorfide):

sorry the coding messed up on that limit with e

OpenStudy (anonymous):

Should the e be after the limit?

OpenStudy (amorfide):

no the limit is in the power

OpenStudy (amorfide):

if it is the negative infinite that you are confused about then

OpenStudy (anonymous):

Yes

OpenStudy (amorfide):

oops some of it was cut off

OpenStudy (anonymous):

Is it because ln(2/9) is negative?

OpenStudy (amorfide):

http://prntscr.com/6ggygk

OpenStudy (amorfide):

yeah

OpenStudy (amorfide):

because it would be like -ln(9/2) or something

OpenStudy (amorfide):

so you get -infinite

OpenStudy (anonymous):

Ok. And in the original thing that you posted here, you can raise the thing to the e power and take the ln of the stuff because they cancel out right?

OpenStudy (anonymous):

http://prntscr.com/6ggqj2

OpenStudy (amorfide):

yeah because e and the ln cancel that means it still turns out to be (2/9)^x but written in another form that is easier to get a limit from

OpenStudy (anonymous):

Ok. I get it now! Thank you so much!!!

OpenStudy (amorfide):

sorry for not being elaborate in the first place lol

OpenStudy (anonymous):

No worries. I don't think that my teacher will make us explain that far, but its hard to tell in this chapter what we have to prove and what we can assume that they know

OpenStudy (amorfide):

I think it is safe to say that aslong as you simplified it to something you can get the limit of such as the e^x format you can just say what the limit is, because you showed the understanding, but you are right, better safe than sorry

OpenStudy (anonymous):

Ok. Thanks again!